OpenStudy (anonymous):

Select the system of equations that corresponds to the given graph. graph of two lines. One line goes through the points 0 comma 2 and 1 comma 0. The other line goes through the points 0 comma 2 and negative 1 comma negative 1 4x + y = 6 6x + 3y = 6 −3x + y = 2 4x + 2y = 4 2x + y = 4 6x + 3y = 12 4x + 2y = 6 6x + 3y = 12

OpenStudy (ddcamp):

Find the 2 equations using the point slope form: \[m = \frac{ y₂ - y₁ }{ x₂ - x₁ }\] y - y₁ = m(x-x₁) Then put the formula in standard form (ax + by = c)

OpenStudy (anonymous):

what would be y and y1 and x and x1